Explore topic-wise InterviewSolutions in Current Affairs.

This section includes 7 InterviewSolutions, each offering curated multiple-choice questions to sharpen your Current Affairs knowledge and support exam preparation. Choose a topic below to get started.

1.

Obtain all other zeroes of 3x raise to power 4 +6x raise to power 3 -2x square-10x-5,if two of its zeroes are √5/3 and - √5/3.BEST ANSWER WILL BE BRAINILIST​

Answer»

hope it HELPS you... THANKS

2.

Дуdx dy dz(D)дхaz3. For the level surface 0(x, y, z) =c, gradient of o represents(A) unive vector(B) tangent vector(C) normal vector(D) radius vector​

Answer»

ANSWER:

XXX

with SUNNY LEONE vghhghiii

3.

5. Prove that the perpendicular at the point of contact to the tangent to a circle passesthrough the centre​

Answer»

Step-by-step EXPLANATION:

Let O be the centre of the given circle. AB is the tangent drawn TOUCHING the circle at A. DRAW AC ⊥ AB at point A, such that point C lies on the given circle. ∠OAB = 90° (Radius of the circle is perpendicular to the tangent) Given ∠CAB  = 90°   ∴ ∠OAB = ∠CAB This is POSSIBLE only when centre O lies on the line AC. Hence, perpendicular at the point of CONTACT to the tangent to a circle passes through the centre of the circle.

4.

Given A=101 o find a maximum setsОО1 - 3 3of Linearly independent eigenivectors OfA.is A diagonalizable? If so find D.​

Answer»

ANSWER:

xxxxxxx

Step-by-step EXPLANATION:

ccffb. vbhfgjb VFG

5.

Plz help me it's urgent this is not on Google and on doutnut also plz help me

Answer»

Step-by-step explanation:

A=40,000

Let part invested in 4 % be x

so principal = x

I(4) = x * 4 * 1/100 = 4x/100

let part invested at 6 % be y

so principal in this case = y

so I(6) = y * 6 * 1/100 = 6y/100

now total I = 2130

which implies (4x + 6y)= 213000

and also A= 40000 = x +y +2130

so x + y = 37870

SOLVING both EQS we GET y = 30760 and x= 7110

hence first amount is 7110 and second is 30760

hope this helps you!!

6.

The probability of an event which is impossible to happen is...​

Answer»

Answer:

0/ No. of POSSIBILITIES = 0

7.

If the square of y varies directly as the cube root of x and y=2,when x=9. Find x when y=4​

Answer»

ANSWER:

Through this process you can SOLVE it.

Step-by-step explanation:

Please mark me as BRAINLIEST ☺️.

8.

Multiplicative inverse of -3/13??????​

Answer»

ANSWER:

-13/3 is the ans.

Solution:

1/-3/13 = 13/-3 = -3/13

9.

The product of 1.2 multiply by 0.6 multiply by 3.12 multiply by 0.03 is equal to ​

Answer»

Answer:

67.452

Step-by-step explanation:

1.2 X 0.6 x 3.12 x 0.03

7.2 x 9.36

67.452

10.

8/5*-3/2-(-3/16*-11/10)​

Answer»

Answer:

Answer is in the above PICTURE

HOPE so it's helpful mark as brainlist...FOLLOW me....❤❤

11.

cross multiplication class 10th the question 3 X + 2 Y equal to 5 and 2 x minus 3 Y equal to 7 please tell me the answer please​ Pls guys help me

Answer»

ANSWER:

Hey mate here is your answer

Step-by-step explanation:

HOPE this may HELP you

Mark me BRAINLEST

12.

What are the steps????? ​

Answer»

Step-by-step EXPLANATION:

tana=5/4

using PYTHAGORAS there.

13.

Cos 21°÷sin69°-sin40°÷cos50°​

Answer»

Step-by-step EXPLANATION:

ANSWER:-

We know that

Sin theta = Cos ( 90- theta)

So is the inverse ie

Cos theta = Sin ( 90- theta)

  • So we can write cos 21° as Sin 69°

And ALSO

  • we can write sin 40° as Cos 50°

If we replace the values we GET

Sin 21°/ Sin 21° - cos 50°/ cos 50°

= 1-1

So the answer is ZERO.

SOME IDENTITIES

  • Sin0 = cos(90-0)
  • Cos0= sin(90-0)
  • Tan0 = cot(90-0)
  • Cot0 = tan(90-0)
  • Sec0 = cosec(90-0)
  • Cosec = sec(90-0)
  • These all are complementary of each other
  • Where 0 = theta
  • sin is Inverse of cosec
  • cos is Inverse of sec
  • tan is Inverse of cot
  • And VICE versa are also correct

14.

Please solve question number 6 and 7 . The answer of question number 6 will be -1/2 and answer number 7 will be A = 45 degree , B = 15 degreePlease solve it fast I have given the question in the image just click it you will get​

Answer»

HEY, user

here is UR SOLUTION...

CHECK out

15.

Please give me the answer with proper written steps. It's very important.

Answer»

Answer:

x=1

FIRST square both sides

APPLY A +B whole square fromula on LHS

in the end you would GET x=1

sorry my answer is too SHORT but i CANT type everything in here

16.

Chapter circleProve that the tangents drawn at the ends of a diameter of a circle are parallel.please answer this question​

Answer»

ANSWER:

hy..your answer in pic above which is attached..

hope it helps

mark brainliest plz...

17.

How can you draw a line with just one point

Answer»

Answer:

???? Can't UNDERSTAND

If we have TWO points A and B such that we can DRAW a line

18.

6. FIND THE VALUE OF P AND Q IF PS LOOK THE QUESTION 6

Answer»

ANSWER:

you can do it by RATIONALISING the DENOMINATOR

19.

For a car the data is as given below,Weight of car is 44 kN, frontal area is 5.4 m2, diameter of wheel is 140 cm,Rolling resistance is 3 N/ 1 kN, constant for air resistance is 0.028, on a flat road car can travel at max. Speed of 140 km/hr. Find out Engine power required to travel at max speed in top gear on flat road.​

Answer»

Step-by-step explanation:

For a car the data is as given below,

WEIGHT of car is 44 kN, frontal area is 5.4 m2

, diameter of wheel is 140 CM,

ROLLING resistance is 3 N/ 1 kN, constant for air resistance is 0.028, on a flat

road car can travel at max. Speed of 140 km/hr. Find out Engine power

required to travel at max speed in TOP gear on flat road.

20.

(cos a -cos3a )(sin8a+sin2a)/(sin5a-sina)(cos4a-cos6a)=?​

Answer»

Step-by-step EXPLANATION:

(COS a -COS3A )(sin8a+sin2a)/(sin5a-sina)(cos4a-cos6a)

(cos a -cos3a )(sin8a+sin2a)/(sin5a-sina)(cos4a-cos6a)=?

21.

Divide p(x)=x⁴+x³+3x+2x+2 by x²+2​

Answer»

Answer:

SORRY I don't understand your QUESTION please MARK me brainlist

22.

L woite all the fractors of the followingWumbeas224 025© 22③ 27 @12209 28 623 36​

Answer»

ANSWER:

QUESTION is INCORRECT.

23.

Multiply 6√7 by 2√7plz solve it I will mark you brilliant ​

Answer»

3...........the answer

6

24.

What is the multiplicative inverse of -3/13​

Answer»

the MULTIPLICATIVE INVERSE of -3/13

=  \frac{ - 13}{3}

hope it will HELP you

25.

Write t the coordinates of the point which​

Answer»

Answer:

Step-by-step explanation:

The COORDINATES of a point are a pair of numbers that define its exact location on a two-dimensional plane. Recall that the coordinate plane has two axes at right angles to each other, called the x and y axis. The coordinates of a given point represent how far along each axis the point is located.

Ordered Pair

The coordinates are WRITTEN as an "ordered pair" as shown below. The letter P is simply the name of the point and is used to distinguish it from others.

Ordered pair SHOWING coordinates of a point The two numbers in parentheses are the x and y coordinate of the point. The first number (x) specifies how far along the x (horizontal) axis the point is. The second is the y coordinate and specifies how far up or down the y axis to go. It is called an ordered pair because the order of the two numbers matters - the first is ALWAYS the x (horizontal) coordinate.

The sign of the coordinate is important. A positive number means to go to the right (x) or up(y). NEGATIVE numbers mean to go left (x) or down (y). (The figure at the top of the page has the values of the axes labelled with the appropriate sign).

26.

Simplify the following surds1)√272)√503)√2504)√1125)√168​

Answer»

Answer:

1)3root3

2)5root2

3)5root10

4)4root7

5)2root42

Step-by-step EXPLANATION:

HOPE it HELPS you

27.

Please answer it fast !!!!

Answer»

Answer:

We have, cosec A = 2

Now, sin A = 

⇒ sin A = 

We KNOW that sin² A + cos² A = 1

⇒ cos² A = 1 - sin² A

⇒ cos A = √(1-sin² A)

⇒ cos A = √{1- }

⇒ cos A = √(1 - )

⇒ cos A = √()

⇒ cos A = 

We know that tan A = 

⇒ tan A = 

⇒ tan A = 

We need to find  + 

Substituting the VALUES of each trigonometric FUNCTION obtained EARLIER, we get:

+  =  + 

=  + 

=  + 

= 

= 

=  × 

=  = 8-6= 2

Step-by-step explanation:

28.

By using Euclid's division algorithm find HCF of305,793

Answer»

Step-by-step explanation:

305,798= 30000050007008

29.

A dozen eggs cost $140. How many eggs can be bought with $250?

Answer»

Answer:

21

Step-by-step EXPLANATION:

30.

Q9. What would be the denominator after rationalizing 7/(5root3 - 5root2 )?

Answer»

ANSWER:

The answer of this QUESTION is 25

31.

Find the product of the successor of 999999 and the predecessor of 1000

Answer»

Answer:

Step-by-step explanation:

SUCCESSOR of 999999 = 1000000

predecessor of 1000= 999

product = 999* 1000000

=999000000

32.

Find the product:(x²-y²)(x²+y²)

Answer»

ANSWER:

x^4 - y ^ 4

Step-by-step EXPLANATION:

this is the answer

33.

In how many ways can we arrange the word fuzztone so that all the vowels come together

Answer»

6 FACTORIAL DIVIDE by 2 factorial and multiply by 3 factorial....

34.

Write the greatest number of different digits formed by using the digits 2,0,8,5

Answer»

4 FACTORIAL =24 is the answer

mark me as BRAINLIEST

35.

Calculate the surface area of a hemispherical dome of a temple with radius 14cm to be white washed from outside. Also find the cost at Rs. 7 per m^2.

Answer»

ANSWER:

SURFACE area of HEMISPHERICAL dome = 2πr^2

= 2×22× 196/7 = 44 × 28 = 1232 sq.cm

= 0.1232 sq.m

total cost = 7 × 0.1232 = 0.8624 /-

hope it helps U !!!!

plzzz mark me as brainliest .....

36.

1.Find the lateral and total surface area of cuboid of length 80cm, 60cm and 45cm.

Answer»

Step-by-step EXPLANATION:

LENGTH(l)=80 cm

BREADTH(B)=40 cm

HEIGHT(h)=20 cm

TSA OF CUBOID = 2(lb+bh+lh)

                             = 2[(80*40)+(40*20)+(20*80)] cm²

                             =2[3200+800+1600] cm²

                             =2[5600] cm²

                             =11200 cm²

LSA OF CUBOID=2(l+b)h

                            =2(80+40)20 cm²

                            =2(120)20 cm²

                            =40(120) cm²

                            =480 cm²

37.

What is the area of a circle that has a diamter of 24 m

Answer»

ANSWER:

Sorry for the scribbling and don't forget to add the unit to the answer as I forgot.

Step-by-step explanation:

If you are satisfied, then PLEASE mark it as a brainiest answer.

HAVE A GOOD DAY AHEAD☺️☺️

38.

The word form of 9000891 in the international system of numeration is

Answer»

nine million EIGHT HUNDRED NINETY ONE

39.

Verify Distributive Property[X×(Y+Z) = X×Y+X×Z] for-X=1/2, Y=-2/3, Z=5/6​

Answer»

ANSWER:

Hii mate here's your answer for this QUESTION...

(1/2×(-2/3+5/6)=1/2×-2/3+1/2×5/6

=> 1/2×-2/3+1/2×5/6=-1/3+5/12

=> -1/3 +5/12 =-1/3+5/12

=> -4+5/12=-4+5/12

=> -1/12 = -1/12

=> LHS=RHS

Step-by-step EXPLANATION:

HOPE this answer helps you mate....

40.

What fraction is the 25 year of 1 century

Answer»

Answer:

25 YEAR of 1 CENTURY

1 century =100 YEARS

25/100 years =1/4

41.

2under root 3+under root is equal to give me answer

Answer»

Answer:

WRITE QUESTION perfectly in BRAINLY way

42.

An employer earns Ruppes 72000 in 3 months Aa) How much does earn in 7 months b)In how many months will he earns Ruppes 360000

Answer»

Step-by-step EXPLANATION:

IF IT SATISFIES YOU, THEN PLEASE MARK IT AS A BRAINLIEST.

Have a NICE day☺️☺️

43.

Solve:-2x+y=5 and 3x÷3y=10 using elimination method

Answer»

y=5/3

x=1

Step-by-step explanation:

2x+y=5 -----eq(1)

3x+3y=10------eq(2)

[MULTIPLYING each equation with the CORRESPONDING no to eliminate one VARIABLE]

(2x+y=5)*3

(3x+3y=10)*2

[after MULTIPLICATION]

6x+3y=15

6x+6y=20

+ - -

________

-3y= -5

y=5/3----eq(3)

PUTTING value of eq(3) in eq(1)

2x+5/3=5

10x+5/3=5

10x+5=5*3

10x+5=15

10x=15-5

10x=10

x= 1

44.

Verify 7x [(+22) + (-9)] = ((7) * 22] + (7×( -9))​

Answer»

Step-by-step EXPLANATION:

this is your ANSWER .PLEASE FOLLOW me and MARK at brailiant.

45.

The length of a rectangular hall is 4 metres less than three times the breadth of the hall. What is the length, if the breadth is 6 metres

Answer»

ANSWER:

  • Length of rectangle = 6m

GIVEN:

  • length of a rectangular hall is 4 metres less than three times the breadth of the hall.
  • Breadth of the rectangle = 6 metres.

TO FIND:

  • Length of rectangle.

SOLUTION:

Breadth of rectangle = 6 m

Length of rectangle is 4 m less then 3 times the breadth of the wall.

Length = 3(6-4) m

= 3(2) m

= 6 m

Length of rectangle = 6m

ADDITIONAL INFORMATION:

If the QUESTION ASK about finding the area of rectangle then :

Formula:.

  • Area of rectangle = (length)(Breadth)

Here:

Length = 6 m

Breadth = 6 m

Area of rectangle = (6*6)

= 36 m²

NOTE:

Some important formulas:

Area of Square = (Side)²

Area of rectangle = Length*Breadth

Area of rhombus = BASE*Altitude

Area of parallelogram = Base*Altitude

Area of TRIANGLE = (Base*Altitude)/2

46.

What is the new method of formation of tables up to 100

Answer»

Step-by-step EXPLANATION:

आयरन अक्षमाला का टमाटर का दुर्गा चरण पिक्चर

47.

Tanya has ten bills in her wallet. She has a total of $40. If she has one more $5 bill than $10 bills, and two more $1 bills than $5 bills, how many of each does she have?

Answer»

Step-by-step EXPLANATION:

gndthyigp ouykkkjrrhaaaaauyyyeeeehnawgyyawytaaaaaaaaaaaytrttttthsyssrlyseheuurjjaaarrretdlllruooooooodddootsssotttjkljneeedjjjkkke

48.

A number is a _____ when it is a product of same

Answer»

┏─━─━─━─━∞◆∞━─━─━─━─┓

✭✮ӇЄƦЄ ƖƧ ƳƠƲƦ ƛƝƧƜЄƦ✮✭

┗─━─━─━─━∞◆∞━─━─━─━─┛

A NUMBER is a MULTIPLE when it is a PRODUCT of same.

┏─━─━─━─━∞◆∞━─━─━─━─┓

✭Hòpè Ìt Hèĺp§.. Fòĺĺòw @Kaniskaa16✭

┗─━─━─━─━∞◆∞━─━─━─━─┛

49.

Two digit natural no. such that sum of its digits is 9

Answer»

Answer:

18,81,27,72,36,63,45,54,90

Step-by-step explanation:

1+8=9

8+1=9

2+7=9

7+2=9

3+6=9

6+3=9

4+5=9

5+4=9

9+0=9

50.

The probability that a bomb hits a target 1/4. Five bombs are aimed at a target. Find the probability that1. 3 bombs hit the target2. At the most 2 bombs hit the target

Answer»

(1) The probability that  3 BOMBS hit the target is 45/128

(2) the probability that  at most 2 bombs hit the target is 167/256

Step-by-step explanation:

Let an event E be

E : Bomb hits the target

Then

Probability of event E

P(E)=\frac{1}{4}

Therefore,

P(\bar E)=1-\frac{1}{4}=\frac{3}{4}

(1) Let event A be

A: 3 bombs hit the target

P(A)=^5C_3(\frac{1}{4})^3(\frac{3}{4})^2

or P(A)=\frac{5\times 4}{2}\times \frac{9}{4^3\times 4^2}

or P(A)=\frac{45}{128}

(2) Let event B be

B: At most 2 bombs hit the target

P(B) = 1 - Probability that 3 bombs hit the target - Probability that 4 bombs hit the target

P(B)=1-\frac{45}{128}-(\frac{1}{4})^4

P(B)=1-(\frac{45}{128}-\frac{1}{256})

P(B)=1-\frac{89}{256}

P(B)=\frac{167}{256}

Hope this answer is helpful.

Know More:

Q: The probability that a bomb hits a target is 0.4. three bombs can destroy a BRIDGE. if in all 5 bombs are thrown find the probability that the bridge will be destroyed.

Click Here: brainly.in/question/1988802

Q: The probability of a bomb hitting a target is 1/2. the least number of bombs that should be dropped to ensure that probability of the target getting bombed is at least 0.9 is:

Click Here: brainly.in/question/1988793